if y varies directly as x, and y=10 when x=7,find y when x=5
Found 3 solutions by josgarithmetic, greenestamps, Edwin McCravy:
Answer by josgarithmetic(39617)
Answer by greenestamps(13198)



Direct variation means y changes by the same factor as x.

In your example, the x value changes from 7 to 5, a factor of 5/7, so the y value changes by that same factor.

ANSWER: 10*(5/7) = 50/7


Answer by Edwin McCravy(20054)

FORMULA for all "varies ___ as" problems

                              product of "directly" or "jointly" variables
"Varies ___ as" variable = k* --------------------------------------------
                                  product of "inversely" variables

[If one of the types of variables "directly", "jointly", or "inversely"
variables are missing, always put 1 for them.]

The "varies ___ as" variable is y, so write:


                   y = k*-----

The only "directly variable" is x, so write it on the top:

                          x
                   y = k*---- 


There is no "inversely variable", so write "1" on the bottom:


                          x
                   y = k*----- 
                          1
or

                   y = k*x


Now we're ready to substitute the numbers for all the letters except the
constant k from this: "y=10 when x=7". Substituting:

                          
                  10 = k*7 
                          
Solve for k:
                          
                10/7 = k
                          

Now go back to this equation:


                   y = k*x



and substitute 10/7 for the constant k.

                   y = (10/7)*x

Now we substitute x = 5:

                   y = (10/7)*(5)

Then finish it up:

                   y = 50/7 = 7 1/7

Edwin